OpenStudy (anonymous):

solve and check. c/3 6 1/7. Solve for c.

OpenStudy (anonymous):

@DanJS plz help.

OpenStudy (anonymous):

A. 2 1/21 B. 2 1/7 C. 18 1/7 D. 18 3/7

OpenStudy (anonymous):

@lizzyrockefeller

OpenStudy (danjs):

i dont understand the problem, there isn't an = sign

OpenStudy (danjs):

can you retype it

OpenStudy (anonymous):

i got her

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

c/3 = 6 1/7

OpenStudy (anonymous):

Please enter an = sign

OpenStudy (anonymous):

thx

OpenStudy (danjs):

first change 6 1/7 to a improper fraction

OpenStudy (danjs):

\[\frac{ 7*6 + 1 }{ 7 }\]

OpenStudy (danjs):

\[\frac{ c }{ 3 } = \frac{ 43 }{ 7 }\] do you know what to do next?

OpenStudy (anonymous):

not realy I'm getting it though.

OpenStudy (danjs):

ok, you can cross multiply, 7*c = 3*43

OpenStudy (anonymous):

K

OpenStudy (danjs):

then , divide by 7 on both sides c = (3*43)/7

OpenStudy (danjs):

then turn that back to a mixed number

OpenStudy (danjs):

129/7 = ....

OpenStudy (danjs):

how many times does 7 go into 129, then remainder/7

OpenStudy (danjs):

you got it all ?

OpenStudy (anonymous):

7 goes into 129... Wait doing long division.

OpenStudy (danjs):